Write an equation for the situation. Kyle's aunt is now three times as old as Kyle. Five years ago, his aunt was four times as old as Kyle. Let k = Kyle's age now.: Write an equation for the situation. Kyle's aunt is now three times as old as Kyle. Five years ago, his aunt was four times as old as Kyle. Let k = Kyle's age now.
Answer by orca(336)

Let k = Kyle's age now.
Then the age of Kyle's aunt NOW is 3k.
The age of Kyle 5 years age is k-5.
The age of Kyle's aunt 5 years ago is 3k-5.
You need to use the following equation to set up an equation:
"Five years ago, his aunt was four times as old as Kyle"
I.e., The age of Kyle's aunt 5 years ago = 4*(The age of Kyle 5 years age)
So
3k-5=4(k-5)